zarkam21 (zarkam21):

Solve the system of equations below and enter your answer as an ordered pair. 4x+6y=-24 -7x+6y=9

OpenStudy (anonymous):

Multiply second equation by -1. 11x = -33 x = -11 Solve for y.

zarkam21 (zarkam21):

Thats incorrect

zarkam21 (zarkam21):

Thats incorrect

zarkam21 (zarkam21):

Thats incorrect

OpenStudy (anonymous):

Sorry, my mistake. 11x = -33, x = -3 y = -2

zarkam21 (zarkam21):

Solve the system of equations below and enter your answer as an ordered pair. 63x-9y=351 -8x+9y=-21

zarkam21 (zarkam21):

Solve the system of equations below and enter your answer as an ordered pair. 63x-9y=351 -8x+9y=-21

OpenStudy (anonymous):

You can deal with smaller numbers by multiplying both sides of the second equation by -1 and then add the LHS to the LHS and the RHS to the RHS of both equations.

OpenStudy (anonymous):

The 6y s subtract out.

zarkam21 (zarkam21):

Sooo? Im kinda confused

zarkam21 (zarkam21):

Sooo? Im kinda confused

zarkam21 (zarkam21):

Sooo? Im kinda confused

zarkam21 (zarkam21):

Sooo? Im kinda confused

zarkam21 (zarkam21):

Sooo? Im kinda confused